Quad Enterprises is considering a new three-year expansion project that requires an initial fixed asset investment of $2.82 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,120,000 in annual sales, with costs of $815,000. If the tax rate is 30 percent, what is the OCF for this project? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g. 1,234,567.)
Answer #1

depreciation per year = 2820000/3 = 940000

operation cash flow = (sales - costs) * (1-tax rate) + depreciation * tax rate

= (2120000 - 815000) * (1-0.3) + 940000 * 0.3

= 11955001195500